if 51.23 g of lithium reacts with 402.79 g of iron(III) sulfate, how many grams of iron would you expect to form?
1
Expert's answer
2019-02-21T03:23:03-0500

6Li +Fe2(SO4)3 = 3Li2SO4 + 2 Fe

n(Li) = m/M = 51.23/6.94 = 7.38 mol

n(Fe2(SO4)3) =m/M =402.79/399.88 = 1.01 mol

Find limiting reactant:

n(Li)/6 = 7.38/6 = 1.23

n(Fe2(SO4)3)/1 = 1.01/1 = 1.01

1.01<1.23 , Fe2(SO4)3 is a limiting reactant

According to equation 1 mol of Fe2(SO4)3 gives 2 mol of Fe

We have 1.01 mol of Fe2(SO4)3 which give x mol of Fe

1/1.01 = 2/x

x = 2.02 mol

n(Fe) = 2.02 mol

m(Fe) = n*M = 2.02* 55.85 = 112.82 g

Answer: 112.82 g